Write the equation of the line through the given points.
28. (1,5),(0,6)
Someone please help me with this equation,any help would be given much gratitude :)

Answer by checkley71(8403)

28) slope=(6-5)/(0-1)=1/-1=-1 (m)
Y=mX+b
5=-1*1+b
5=-1+b
b=5+1
b=6 THE Y INTERCEPT.
Y=-X+6 IS THE LINE EQUATION.